solve for the structure

The formula is C,HieO 3 H 6H I H The formula is C,Hio to 6 ppm4


solve for the structure
Answer #1

C5H10O2;

Double bond equivalent (DBE) = 5 + 1 - 10/2 = 1

One double bond

C5H10

Double bond equivalent (DBE) = 5 + 1 - 10/2 = 1

One ring

C3H6O

Double bond equivalent (DBE) = 3 + 1 - 6/2 = 1

One double bond

Structures are as follows:

Molecular Formula C5H1002 Molecular Formula CSH10 Molecular Formula C3H6O O 13 5.1 2.1 1.3 1.7 1.7 2.2 2.2